What volume (in mL ) of a 0.175 M HNO 3 solution is required to completely react with 28.6 mL of a 0.108 M Na 2 CO 3 solution according to the following balanced chemical equation? N a 2 C O 3 (aq)+2HN O 3 (aq)→2NaN O 3 (aq)+C O 2 (g)+ H 2 O(l)
Answer #1

Mmol of Na2CO3 = M × V

Mmol = 0.108×28.6 => 3.0888 mmol

From balanced chemical reaction,

MMoles of Na2CO3 =2× mmoles of HNO3

Mmol of HNO3 =2× 3.0888 mmol

Volume(in mL) of HNO3 = MMole/Molarity

V = 2×3.0888/0.175

=>35.3 mL